# Talking to DINO: Bridging Self-Supervised Vision Backbones with Language for Open-Vocabulary Segmentation

### v1

@@ Appendix A Additional Experiments and Analyses
These observations suggest that the global objective of CLIP’s training, similar to its effect on visual patch embeddings, may not endow text tokens with strong local properties that accurately reflect the specific word each embedding represents. This limitation likely contributes to the noisiness of such alignments. Additionally, we evaluate the use of the average of CLIP text tokens in both training and inference as an alternative to the CLS token. While this approach slightly improves over aligning individual tokens, it still underperforms compared to the CLS token, indicating that it encapsulates the most useful and less noisy information for alignment with DINOv2 patches.
mIoU
VOC Context Stuff Cityscapes ADE
Effect of Training CLIP Last Layer
Trained 83.0 36.9 25.8 30.3 19.0
Frozen 85.3 40.5 27.9 38.2 21.8
Effect of Text Token Selection
Average text tokens 83.2 37.8 25.5 32.7 19.6
Text token to best self-attn map 84.2 34.8 24.2 32.9 18.7
Text token to best self-attn map (NS) 81.7 33.4 23.8 29.1 18.8
CLS token only 85.3 40.5 27.9 38.2 21.8
@@
Talk2DINO 32.9 59.2 71.1 3 14.6 12.2 34.3 48.7 11 34.2
   + Custom Alignment 39.4 67.7 78.2 2 10.7 26.2 53.0 66.4 5 21.0
@@ Appendix B Image-Text Matching Results
In Table 8, we assess the retrieval performance on the COCO Captions test set [27] using both ViT-B and ViT-L configurations. While Talk2DINO generally performs slightly below CLIP across most metrics, it demonstrates a notable advantage in the mean rank for the text-to-image retrieval task. This result underscores the ability of Talk2DINO to better address extreme failures compared to CLIP, indicating improved robustness in handling challenging or outlier queries. In addition to computing text-image similarities using cosine similarity between a global text token and a global image token, we experiment with a similarity function that mirrors the one used during training. Specifically, instead of representing the image with the mean of the $v_{A_{i}}$ embeddings and calculating similarity as the cosine similarity between this representation and the text encoding, we represent the image using all $v_{A_{i}}$ embeddings. We compute the similarity as $\max_{i=1,\dots,N}\text{sim}(A_{i},v,t)$, taking the maximum similarity score across all heads. This alternative similarity function leads to significant performance improvements, allowing Talk2DINO to surpass CLIP on several metrics. This enhancement is likely due to the ability of the model to evaluate captions at a finer granularity. Captions often describe multiple aspects of an image, including both foreground and background elements. By individually examining different regions of the image as detected by distinct attention heads, the model can assign more precise scores, ultimately boosting retrieval accuracy.
@@ Appendix C Activation Map Visualizations
In Fig. 6, we show the distribution of attention heads selected for alignment with the text input during the final epoch of training. The results indicate that certain heads, particularly heads 0, 4 (by far the most frequently selected), 8, and 12, are more often aligned with the text than others. However, aside from these, the remaining heads are relatively evenly distributed. These findings are noteworthy because they suggest that some heads specialize in capturing features that align more closely with the input caption, while all heads contribute meaningfully during training. Notably, no head shows a negligible activation frequency, highlighting the importance of the entire set of attention heads in the alignment process.
@@
Fig. 7 presents examples from the training set, showcasing images paired with their corresponding captions and the attention maps selected for alignment. Despite describing the same scene, variations in the captions lead the alignment procedure to focus on different regions of the image. For instance, in the first row, a caption mentioning only the food on the table excludes the dog from the focus, while captions that also reference the dog include it in the relevant patches.
Head 3: A cut up sandwich on top of a wooden table.
Head 8: a dog is stepping up onto the table to eat a plate of food
Head 11: A small dog is sitting at the kitchen table in front of a sandwich.
Head 12: A dog in a shirt sits at a table with a plate of food in front of the dog.
Head 15: A Chihuahua sits at a table next to a sandwich.
Head 0 A baseball games is underway with someone on second base as the pitcher throws the ball.
Head 2: A baseball game from a distance during a pitch.
Head 6: A professional baseball game between to MLB teams
Head 11: baseball players on a baseball field during a professional game
Head 12: Baseball players are ready for the next pitch.
Head 2: A white round table filled with some assorted treats.
Head 5: a dessert tray with dougnuts, cupcakes and muffins
Head 6: A white plate topped with chocolates, donuts and other goodies.
Head 9: A number of pastry items on a table
Head 15: Donuts and other goodies on a table
Head 0: A surfer in a bright green suit expertly takes a wave.
Head 2: a surfer is out at sea riding a wave
Head 4: A surfer turns to catch the oncoming wave.
Head 10: A man who is on a surfboard riding a wave.
Head 14: A man surfing in strong waves in the sea
Head 1: A zebra is standing in an enclosure in a zoo.
Head 3: A zebra is in an enclosure by a fence.
Head 7: A zebra standing near another zebra in a enclosure.
Head 9: A zebra standing in the dirt in an enclosure.
Head 10: A picture of zebra in a enclosed area.
@@ Appendix D Additional Qualitative Results
Effect of Background Cleaning. In Fig. 10 we showcase a set of qualitative results in which we highlight the advantages of using the proposed background cleaning procedure with respect to directly thresholding the similarities with the provided categories to detect the background. In particular, the first two rows show four qualitatives on images from COCO Object and the last two rows from Pascal VOC. These results demonstrate that background cleaning removes the noise in the background from the image and improves the fitting of the masks on the foreground objects. These findings are reflected in the results reported in Table 5 of the main paper.
In-the-Wild Qualitative Examples. Fig. 11 depicts a few examples of “in-the-wild” segmentation, obtained by providing to Talk2DINO some images retrieved from the web and asking it to segment uncommon categories, such as “pikachu”, “millennium falcon”, and “westminster abbey”, and free-form text, like “golden retriever puppy”. On the left, we show three examples in which we task the model with also finding the background, while exploiting the background cleaning procedure, and, on the right, three examples in which the model has to assign a provided category to each pixel. The high quality of the resulting masks demonstrates the efficacy of our approach, even on out-of-domain images with respect to the standard benchmarks. From these examples, we can appreciate the model’s capabilities in combining the knowledge from CLIP with the semantic localization of DINOv2 on unconventional concepts, such as fictional character names and proper nouns of buildings and monuments.
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption 2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
moai dinosaur man westminster abbey golden retriever puppy grass leaves pikachu traffic sign forest street millennium falcon planet space

### v3

@@ Appendix A Additional Experiments and Analyses
These observations suggest that the global objective of the training of CLIP, similar to its effect on visual patch embeddings, may not endow text tokens with strong local properties that accurately reflect the specific word each embedding represents. This limitation likely contributes to the noisiness of such alignments. Additionally, we evaluate the use of the average of CLIP text tokens in both training and inference as an alternative to the CLS token. While this approach slightly improves over aligning individual tokens, it still underperforms compared to the CLS token, indicating that it encapsulates the most useful and less noisy information for alignment with DINOv2 patches.
ViT-Base (mIoU) ViT-Large (mIoU)
Model Visual Backbone Resolution V20 C59 Stuff City ADE Avg V20 C59 Stuff City ADE Avg
without Mask Refinement
SCLIP [47] CLIP 336 80.4 34.2 22.4 32.2 16.1 37.1 70.6 25.2 17.6 21.3 10.9 29.1
NACLIP [18] CLIP 336 79.7 35.2 23.3 35.5 17.4 38.2 78.7 32.1 21.4 31.4 17.3 36.2
ProxyCLIP [26] CLIP+DINOv2 336 80.5 37.3 25.3 35.8 19.0 39.6 83.5 36.7 25.0 35.8 21.0 40.4
ProxyCLIP [26] CLIP+DINO 336 78.2 38.8 26.2 39.7 19.7 40.5 82.1 38.2 26.2 41.2 22.2 42.0
Talk2DINO (Ours) DINOv2 336 88.3 39.1 27.4 38.2 20.2 42.6 86.6 38.2 26.0 36.4 19.3 41.3
with Mask Refinement
SCLIP [47] CLIP 336 79.3 34.6 22.3 20.3 15.4 34.4 66.6 22.4 14.7 6.9 7.7 23.7
NACLIP [18] CLIP 336 83.0 38.4 25.7 38.3 19.1 40.9 84.5 36.4 24.6 37.1 19.6 40.4
ProxyCLIP [26] CLIP+DINOv2 336 80.9 39.3 26.6 37.7 19.9 40.9 83.5 36.7 26.4 38.6 22.1 41.5
ProxyCLIP [26] CLIP+DINO 336 78.5 39.3 26.7 40.1 20.0 40.9 82.6 38.7 26.7 42.1 22.5 42.5
Talk2DINO (Ours) DINOv2 336 89.4 41.5 29.4 40.3 21.2 44.4 89.5 41.7 29.8 38.7 20.8 44.1
without Mask Refinement
SCLIP [47] CLIP 448 77.8 33.0 21.1 19.8 14.6 33.3 61.2 20.5 13.1 6.7 7.0 21.7
NACLIP [18] CLIP 448 71.3 34.8 22.9 33.7 17.7 36.1 74.5 32.6 21.6 30.5 17.8 35.4
ProxyCLIP [26] CLIP+DINOv2 448 83.3 37.8 25.6 28.8 19.1 38.9 85.0 36.6 25.0 33.8 20.6 40.2
ProxyCLIP [26] CLIP+DINO 448 80.4 39.0 26.2 31.7 19.5 39.4 83.1 37.8 25.9 37.5 21.6 41.2
Talk2DINO (Ours) DINOv2 448 87.1 39.8 28.1 36.6 21.1 42.5 87.1 39.1 27.0 35.8 21.1 42.0
with Mask Refinement
SCLIP [47] CLIP 448 79.3 34.6 22.3 20.3 15.4 34.4 66.6 22.4 14.7 6.9 7.7 23.7
NACLIP [18] CLIP 448 74.9 37.6 25.2 36.1 18.4 38.4 79.8 36.8 25.0 35.6 18.4 39.1
ProxyCLIP [26] CLIP+DINOv2 448 83.1 39.3 26.7 29.5 19.7 39.7 85.1 37.8 25.9 35.3 21.4 41.1
ProxyCLIP [26] CLIP+DINO 448 80.0 39.1 26.5 31.7 19.5 39.4 82.8 37.8 26.2 26.2 21.6 38.9
Talk2DINO (Ours) DINOv2 448 88.5 42.4 30.2 38.1 22.5 44.3 89.8 42.7 29.6 38.4 22.9 44.7
Table 7: Comparison with unsupervised OVS models on Pascal VOC [15], Pascal Context [34], COCO Stuff [7], Cityscapes [11], and ADE20K [61, 62] following the evaluation setting proposed in SCLIP [47] (resolution 336) and TCL [9] (resolution 448).
Impact of Image Resolution. According to the evaluation protocol introduced in GroupViT [54] and standardized in TCL [9], the images are resized to have a shorter side of 448, and a sliding window approach with a stride of 224 pixels is employed. However, Wang et al. [47] observed that the approaches based on CLIP benefit from employing a shorter side of 336 with 224 $\times$ 224 windows and a stride of 112 pixels, leading to an equivalent computational effort but better performance. This phenomenon is attributed to two reasons: (i) each window has the same resolution on which CLIP has been originally trained, and (ii) CLIP presents an impressive global understanding but lacks localization capabilities, hence relying on many smaller windows is more advantageous than more patches. This variation of the evaluation setting is not necessary for Talk2DINO because DINOv2, which is the frozen underlying visual encoder, has been trained with a 518 $\times$ 518 resolution and presents an outstanding patch-level understanding, However, Tab. 7 reports the results obtained following the setting used in SCLIP, employing a shorter side of 336 for VOC, Context, COCO-Stuff and ADE, of 560 for Cityscapes, with 224 $\times$ 224 windows and stride 112. Results show that Talk2DINO, on average, performs better with a resolution of 448, but the performance slightly varies when changing the setting to 336. This confirms that the semantics of the patch-level features of DINOv2 are robust towards variations of resolution and that our learned bridge is valid for both scenarios. Moreover, for a fair comparison, we also report the results of SCLIP, NACLIP, and ProxyCLIP when adopting the 448 resolution of the standard protocol, in which Talk2DINO largely outperforms the competitors.
Comparison with ProxyCLIP and FreeDA. GroupViT [54] has been the first model to tackle the weakly-supervised OVS. It trains a custom ViT architecture from scratch by hierarchically merging tokens at different layers. Afterward, several works followed this direction, investigating how to let the segmentation capabilities to emerge by training over a large corpora of image-caption pairs. On the contrary, more recent works focused on finding modifications to the architecture of CLIP in order to improve its localization properties. Moreover, some methods consider the usage of further visual encoders with enhanced localization capabilities to help CLIP on dense tasks. Among these methods, ProxyCLIP and FreeDA study how to combine DINO and DINOv2 with CLIP.
FreeDA employs Stable Diffusion to create a huge collection of localized images from captions, detecting the area in which each noun of the caption has been generated. This information is used to build a database of textual-visual embedding pairs, in which the textual embedding is obtained with CLIP on each noun and the visual embedding is the average patch-level embedding of DINOv2 from the corresponding area. Then, at inference time, a set of textual embeddings is retrieved for each input category, and the corresponding visual embeddings are averaged to create a prototype for that category in the space of DINOv2. Finally, the CLIP visual encoder runs on the input image to solve ambiguities and remove noise. ProxyCLIP proposes to leverage the semantic coherence of a visual encoder such as DINO or DINOv2 to guide the computation of the patch-level embeddings of CLIP. This guidance is performed inside an attention module, in which the patch-level embeddings of DINO act as queries and keys while those of CLIP act as values.
@@
Talk2DINO 29.5 56.0 69.0 4 16.4 12.5 34.0 48.4 11 38.4
  + Custom Alignment 28.6 58.8 72.0 4 12.0 28.0 55.6 68.7 4 19.3
@@
Visual Encoder Params (M) FLOPS (G) Ext. (GiB)
ProxyCLIP CLIP ViT-B/16 + DINO ViT-B/8 172.0 521.2 -
ProxyCLIP CLIP ViT-B/16 + DINOv2 ViT-B/14 172.8 180.8 -
FreeDA CLIP ViT-B/16 + DINOv2 ViT-B/14 172.8 125.1 12.5
Talk2DINO DINOv2 ViT-B/14 86.6 107.4 -
Table 9: Number of parameters, FLOPS, and the dimension of the external knowledge for ProxyCLIP, FreeDA, and Talk2DINO.
Model Visual Encoder V20 C59 Stuff City ADE V21 C60 Object Avg
DINOv2 ViT-B/14 with registers (without Mask Refinement)
FreeDA DINOv2 83.4 39.5 25.9 35.2 20.7 50.1 $\triangleright$ 43.6 34.3 23.8 $\triangleright$ 24.7 39.1 $\triangleright$ 38.4
FreeDA CLIP+DINOv2 87.0 40.6 25.7 34.2 21.2 49.3 $\triangleright$ 41.8 35.7 34.8 $\triangleright$ 34.7 41.1 $\triangleright$ 40.1
ProxyCLIP CLIP+DINOv2 83.0 37.2 25.4 33.9 19.7 58.6 $\triangleright$ 60.0 33.8 37.4 $\triangleright$ 37.3 41.1 $\triangleright$ 41.3
Talk2DINO DINOv2 87.1 39.8 28.1 39.6 21.1 59.9 $\triangleright$ 61.5 35.1 37.1 $\triangleright$ 41.0 43.5 $\triangleright$ 44.2
DINOv2 ViT-B/14 with registers (with Mask Refinement)
FreeDA DINOv2 84.9 42.3 27.7 36.8 22.0 50.2 $\triangleright$ 43.7 36.7 24.5 $\triangleright$ 25.5 40.6 $\triangleright$ 40.0
FreeDA CLIP+DINOv2 87.4 42.4 26.6 34.8 22.1 49.4 $\triangleright$ 41.7 37.2 36.6 $\triangleright$ 36.7 42.1 $\triangleright$ 41.1
ProxyCLIP CLIP+DINOv2 83.1 38.9 26.6 35.4 20.3 62.0 $\triangleright$ 63.4 35.2 38.7 $\triangleright$ 38.6 42.5 $\triangleright$ 42.7
Talk2DINO DINOv2 88.5 42.4 30.2 41.6 22.5 63.9 $\triangleright$ 65.8 37.7 40.3 $\triangleright$ 45.1 45.9 $\triangleright$ 46.7
Table 10: Comparison between FreeDA, ProxyCLIP, and Talk2DINO when using DINOv2 with and without registers. For VOC21, Object, and the average, we report the results without background cleaning on the left and with background cleaning on the right.
Talk2DINO, similarly to FreeDA and ProxyCLIP, investigates how to leverage DINOv2 to compensate for CLIP. However, we propose to employ contrastive learning over a large set of image-caption pairs based on maximum similarity between the attention head embeddings and texts, to learn a functional mapping that bridges the CLIP text embeddings into the DINOv2 space. Our approach demonstrates that the two spaces can be directly connected to set the new state-of-the-art in the unsupervised OVS field. Table 9 shows a quantitative comparison in terms of the number of parameters and FLOPS of the visual encoders and the dimension of the external knowledge (i.e., the database of FreeDA), when assuming an input image with a resolution of 448 $\times$ 448. The results highlight that our method is more practical and less demanding in computation and memory, while presenting improved results against all competitors.
In Tab. 1, we followed the original configurations of the competitors and, hence, ProxyCLIP uses DINOv2 with registers while FreeDA does not. We report a comparison with and without registers in Tab. 10. The registers present the greatest impact on Talk2DINO, because, as described in ”Role of DINO Registers”, the presence of anomaly tokens leads all the self-attention heads to focus only on them, preventing the selection of diverse areas during training and, hence, limiting the efficacy of our proposal. Moreover, in Tab. 10 we report the effect of the background cleaning also on FreeDA and ProxyCLIP. This approach is effective only on Talk2DINO due to the learned alignment between text and average embeddings of the self-attention heads, while it leads to lower results when applied to the other methods.
ViT-B vs ViT-L. Tab. 1 of the main paper shows that, without mask refinement, the results achieved by Talk2DINO with DINOv2 ViT-B as vision encoder are slightly better than the ones achieved with ViT-L, while the opposite should be expected. However, when we apply the PAMR for mask refinement, the results of ViT-L significantly improve, surpassing the ViT-B on five benchmarks out of eight. A similar phenomenon can be observed in other competitors, such as MaskCLIP, SCLIP, ClearCLIP, and NACLIP, while in FreeDA and ProxyCLIP we cannot establish an encoder size that prevails on the other. Even from the experiment in Tab. 5 on patch-level linear probing, we can observe that ViT-B performs slightly better than ViT-L. These results suggest that DINOv2 ViT-L has a comparable semantic understanding with respect to ViT-B, but presents inferior localization properties, which are compensated through PAMR. We hypothesize that training the model with a form of weak- or self-supervision by exploiting the innate capabilities of pre-trained backbones lacks a direct relation between performance and model size. Indeed, the impressive semantic and localized understanding of DINOv2 is a consequence of its training procedure but not the direct objective. From Figure 8, it is noteworthy that the activations of ViT-S, ViT-B, and ViT-L have very different behaviors, impacting the results of Talk2DINO.
@@ Appendix B Image-Text Matching Results
In Table 8, we assess the retrieval performance on the COCO Captions test set [31] using both ViT-B and ViT-L configurations. While Talk2DINO generally performs slightly below CLIP across most metrics, it demonstrates a notable advantage in the mean rank for the text-to-image retrieval task. This result underscores the ability of Talk2DINO to better address extreme failures compared to CLIP, indicating improved robustness in handling challenging or outlier queries. In addition to computing text-image similarities using cosine similarity between a global text token and a global image token, we experiment with a similarity function that mirrors the one used during training. Specifically, instead of representing the image with the mean of the $v^{A_{i}}$ embeddings and calculating similarity as the cosine similarity between this representation and the text encoding, we represent the image using all $v^{A_{i}}$ embeddings. We compute the similarity as $\max_{i=1,\dots,N}\text{sim}(v^{A_{i}},t)$, taking the maximum similarity score across all heads. This alternative similarity function leads to significant performance improvements, allowing Talk2DINO to surpass CLIP on several metrics. This enhancement is likely due to the ability of the model to evaluate captions at a finer granularity. Captions often describe multiple aspects of an image, including both foreground and background elements. By individually examining different regions of the image as detected by distinct attention heads, the model can assign more precise scores, ultimately boosting retrieval accuracy.
@@ Appendix C Activation Map Visualizations
In Fig. 6, we show the distribution of attention heads selected for alignment with the text input during the final epoch of training. The results indicate that certain heads, particularly heads 1 and 3, are more often aligned with the text than others. However, aside from these, the remaining heads are relatively evenly distributed. These findings are noteworthy because they suggest that some heads specialize in capturing features that align more closely with the input caption, while all heads contribute meaningfully during training. Notably, no head shows a negligible activation frequency, highlighting the importance of the entire set of attention heads in the alignment process.
@@
Fig. 7 presents examples from the training set, showcasing images paired with their corresponding captions and the attention maps selected for alignment. Despite describing the same scene, variations in the captions lead the alignment procedure to focus on different regions of the image. For instance, in the first row, the caption mentioning the fans also focuses on the background, while captions that reference only the player, the ball, and the racket do not.
Head 1: A professional tennis player hits a ball as fans watch.
Head 5: champion tennis player swats at the ball hoping to win
Head 7: A man hitting a tennis ball with a racquet.
Head 10: A man is hitting his tennis ball with a recket on the court.
Head 11: a tennis player on a court with a racket
Head 0: A gray and white building on the corner of bay street with building behind it.
Head 1: A street in the ciy San franscisco on a good day
Head 2: A tree leaning on a building on Bay Street.
Head 5: A car driving down a street next to a tall building.
Head 9: A street sign for Bay Street in a residential neighborhood.
Head 1: A puppy holding a black disk on the couch
Head 2: A dog sitting on a couch holding a frisbee in its mouth.
Head 3: A dog holding a plate while sitting on a chair.
Head 9: A bull dog holding a frisbee in it’s mouth.
Head 10: a brown black and white dog and a black frisbee
Head 0: several of the baseball players are in view
Head 1: A large crowd is watching a baseball game.
Head 3: Players on a baseball field during a game.
Head 5: A man is up to bat at a professional baseball game.
Head 6: a man holding up his baseball bat during a baseball game
Head 1: A white round table filled with some assorted treats.
Head 6: A number of pastry items on a table
Head 7: A dessert tray with donuts, cupcakes, and muffins
Head 10: A white plate topped with chocolates, donuts, and other goodies.
Head 11: Donuts and other goodies on a table
@@ Appendix D Additional Qualitative Results
Effect of Background Cleaning. Fig. 10 shows a set of qualitative results in which we highlight the advantages of using the proposed background cleaning procedure with respect to directly thresholding the similarities with the input categories to detect the background. In particular, the first two rows show four qualitatives on images from COCO Object and the last two rows from VOC. These results demonstrate that background cleaning removes the noise in the background from the image and improves the fitting of the masks on the foreground objects. These findings are reflected in the results reported in Table 4 of the main paper.
In-the-Wild Qualitative Examples. Fig. 11 depicts a few examples of “in-the-wild” segmentation, obtained by providing to Talk2DINO sample images from the web and asking it to segment uncommon categories, such as “pikachu”, “millennium falcon”, and “westminster abbey”, and free-form text, like “golden retriever puppy”. On the left, we show three examples in which we task the model with also finding the background, while exploiting the background cleaning procedure, and, on the right, three examples in which the model has to assign a provided category to each pixel. The high quality of the resulting masks demonstrates the efficacy of our approach, even on out-of-domain images. From these examples, we can appreciate the capabilities of the model in combining the knowledge from CLIP with the semantic localization of DINOv2 on unconventional concepts, such as fictional character names and proper nouns of historical buildings.
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption 2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
millennium falcon tordelli westminster abbey golden retriever puppy grass field vegetation pikachu traffic sign forest route axolotl stones leaves gravel